# Protein Design with Dynamic Protein Vocabulary

### v1

@@ 1 Introduction
Figure 1: (a) Visualization of proteins designed by our method, embedded using ESM C hayes2024simulating and projected with UMAP mcinnes2018umap . Random refers to proteins generated by randomly selecting amino acids according to their empirical distribution in SwissProt bairoch2000swiss . Random+ refers to proteins generated by selecting amino acids and incorporating fragments. Natural proteins randomly sampled from SwissProt (in gray) form a broad distribution, representing the diverse landscape of natural proteins. Proteins generated by Random sampling from all possible sequences (in red scatter points) cluster tightly at the periphery of the natural protein distribution, suggesting that random proteins are less diverse. Proteins generated by Random+ (in yellow scatter points) exhibit a much more diverse distribution than those generated by Random. (b) Performance on pLDDT. (c) Performance on PAE.
@@ 2 Background
Proteins are macromolecules composed of linear chains of amino acids, with 20 standard amino acids serving as their fundamental building blocks. Continuous subsequences $S=\{s_{1},s_{2},\cdots,s_{m}\}$ within an amino acid sequence $P$, often referred to as fragments or segments, play crucial roles in determining the structure and function of proteins. In Figure 2, we utilize InterPro blum2025interpro , an integrated database and diagnostic tool, to identify these fragments within proteins. These fragments can be classified into 7 categories: Domain, Family, Homologous Superfamily, Repeat, Conserved Site, Active Site, Binding Site, and Post-Translational Modification (PTM). Details of fragment types are discussed in Appendix B. The fragments, along with their corresponding types and descriptions, are referred to as functional annotations, denoted by $\mathcal{A}_{\text{type}}(S)$ and $\mathcal{A}_{\text{desc}}(S)$.
@@ 3.2 Training
$\mathbf{W}_{\text{in}}&=\mathrm{Concat}(\mathbf{W}_{\text{tokens, in}},\mathbf{W}_{\text{fragments}})\\ \mathbf{H}_{\text{pre}}&=\mathrm{Concat}(\mathbf{H}_{t},\mathbf{W}_{\text{in}}[x_{<i}])\end{aligned}\right.$ (1)
@@
$\mathbf{W}_{\text{out}}&=\mathrm{Concat}(\mathbf{W}_{\text{tokens, out}},\mathbf{W}_{\text{fragments}}^{\mathrm{T}})\\ p(x_{i}=k|\mathbf{H}_{\text{pre}})&=\frac{\exp(\mathbf{H}_{\text{pre}}\mathbf{W}_{\text{out}}^{(k)})}{\sum\limits_{k^{\prime}\in V_{\text{tokens}}\cup S}\exp(\mathbf{H}_{\text{pre}}\mathbf{W}_{\text{out}}^{(k^{\prime})})}\end{aligned}\right.$ (2)
@@ 4.1 Experimental Setups
To comprehensively evaluate our approach, we consider the following metrics from three perspectives. Detailed definitions and implementations are provided in Appendix D. Sequence Plausibility. For rapid validation of designed protein sequences, we employ Perplexity (PPL) and Repetitiveness (Rep) welleck2019neural to evaluate sequence-level plausibility. Foldability. Following dai2024toward , we utilize ESMFold lin2022language to predict the 3D structures of designed proteins and compute the predicted Local Distance Difference Test (pLDDT) jumper2021highly as a per-residue measure of local confidence (0-100, higher is better). Additionally, we also compute the Predicted Aligned Error (PAE) jumper2021highly , which measures the confidence in the relative position of two residues within the predicted structure (lower is better). Consistent with previous researches jumper2021highly ; akdel2022structural ; varadi2022alphafold ; dai2024toward , we adopt pLDDT above 70 and PAE below 10 to define well-folded predicted structures. Language Alignment. To mitigate the high cost of wet-lab experiments, we employ oracle model-based metrics to evaluate the alignment between proteins and functions. Following dai2024toward , we use ProTrek su2024protrek , which evaluates alignment using cosine similarity, referred to as the ProTrek Score, within a joint embedding space. We also use EvoLlama liu2024evollama , fine-tuned on our downstream datasets, computing the EvoLlama Score as the cosine similarity between textual descriptions and predicted functions. Higher scores from both models indicate better alignment with the specified function. Additionally, we employ two retrieval-based metrics. The first is Keyword Recovery hayes2024simulating , which measures the recovery of function keywords in the designed proteins using InterProScan jones2014interproscan . And the second is Retrieval Accuracy liu2025text , which assesses how well the designed proteins aligns with their functional descriptions by evaluating whether their representations exhibit the highest similarity among all candidate pairs.
@@ 4.2 Designing Proteins from Function Keywords
Models #Pairs #Params Sequence Plausibility Foldability Language Alignment (in %)
PPL ($\downarrow$) Rep ($\downarrow$) pLDDT ($\uparrow$) % > 70 ($\uparrow$) PAE ($\downarrow$) % < 10 ($\uparrow$) ProTrek Score ($\uparrow$) Keyword Recovery ($\uparrow$) Retrieval Accuracy@20 ($\uparrow$)
Natural - 467.64 0.02 81.21 90.53 7.08 82.05 21.09 100.00 70.81
Random (U) - 2471.95 0.01 24.38 0.00 23.81 0.13 7.50 0.00 6.05
Random (E) - 3046.64 0.01 27.46 0.00 23.70 0.00 6.59 0.00 5.13
Random+ (E) - 966.24 0.01 62.38 32.65 17.23 9.28 3.29 0.00 5.79
ProteinDT 541K/729M 1405.70 0.11 38.70 0.20 26.25 0.00 3.89 0.05 7.43
ProteinDTFT 392K/729M 1860.43 0.04 38.66 1.04 23.90 0.42 6.28 1.08 16.57
Pinal 1.76B/2B 584.22 0.15 66.50 47.21 14.57 33.53 14.57 30.46 51.68
PAAG 130K/1.3B 2571.40 0.02 33.14 0.00 23.31 0.00 5.21 0.23 7.10
PAAGFT 392K/1.3B 2004.01 0.04 41.53 1.12 24.34 0.46 3.46 0.01 7.82
Chroma 45K/334M 1322.37 0.03 61.66 28.96 13.01 39.03 2.97 0.11 6.57
ESM3 539M/1.4B 279.78 0.33 59.79 31.49 17.40 21.37 3.76 5.49 11.97
ProDVa 392K/1.8B 656.04 0.01 75.88 77.00 6.39 83.88 14.43 30.34 44.77
@@ 4.3 Designing Proteins from Textual Descriptions
Models #Pairs #Params Sequence Plausibility Foldability Language Alignment (in %)
PPL ($\downarrow$) Rep ($\downarrow$) pLDDT ($\uparrow$) % > 70 ($\uparrow$) PAE ($\downarrow$) % < 10 ($\uparrow$) ProTrek Score ($\uparrow$) EvoLlama Score ($\uparrow$) Retrieval Accuracy@20 ($\uparrow$)
Natural - 318.15 0.02 80.64 81.27 9.20 65.73 27.00 60.33 84.85
Random (U) - 2484.03 0.01 22.96 0.16 24.85 0.56 1.03 36.23 6.89
Random (E) - 3136.88 0.01 25.77 0.20 24.71 0.60 1.04 34.11 6.78
Random+ (E) - 846.01 0.01 64.47 37.03 17.91 7.52 0.30 38.65 6.13
ProteinDT 541K/729M 1576.23 0.07 38.29 0.98 25.13 0.40 1.20 40.57 9.28
ProteinDTFT 712K/729M 1213.38 0.04 51.42 25.61 18.57 23.92 13.89 52.84 47.29
Pinal 1.76B/2B 308.97 0.13 75.25 68.97 10.96 58.44 17.50 53.42 57.95
PAAG 130K/1.3B 2782.70 0.02 28.39 0.07 25.38 0.10 1.29 34.39 7.06
PAAGFT 712K/1.3B 1332.35 0.04 50.37 23.86 19.96 21.99 10.04 49.69 33.66
Chroma 45K/334M 1370.21 0.03 59.18 20.17 15.03 28.62 2.10 40.10 7.33
ProDVa 712K/1.8B 415.63 0.02 76.86 76.35 8.66 68.06 17.40 51.10 59.07
@@ Appendix B Details of InterPro and Fragment Types
Table 3 presents the details of the 7 fragment types identified by InterPro.
Table 3: Descriptions of 7 fragment types identified by InterPro.
@@ C.1 Training Details
ProDVa is trained using the AdamW optimizer [52] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a gradient clipping of 1.0. The maximum learning rate is set to $1\times 10^{-4}$, with a linear warmup over the first 5% of training steps. The learning rate remains constant after warmup and follows a "1-sqrt" decay schedule during the last 10% of training steps. The minibatch size is set to 4 to ensure memory efficiency, with an overall batch size of 64. Training is conducted for 10K steps on the CAMEO subset and 20K steps on Mol-Instructions. The loss weights are set to $\alpha=\beta=0.2$.
@@ D.1 Sequence Plausibility Metrics
where $(x_{1},x_{2},\cdots,x_{n})$ is a sequence of tokens tokenized using the ProtGPT2 tokenizer.
@@ D.3 Language Alignment Metrics
The ProTrek Score is defined as the cosine similarity between a textual function description $t$ and a designed protein sequence $P$. Since ProDVa generates protein sequences directly, without an intermediate step for designing structure backbones, only the text and sequence encoders of ProTrek [32] are utilized. Specifically, we employ the pre-trained ProTrek 650M  [^3] https://huggingface.co/westlake-repl/ProTrek_650M_UniRef50 as the oracle model. The ProTrek Score is formally defined as:
ProTrek $\text{ProTrek Score}=\mathrm{sim}(\tau_{t}(t),\tau_{p}(P))$
where $\mathrm{sim}(\cdot,\cdot)$ denotes the cosine similarity function, and $\tau_{t}(t)$ and $\tau_{p}(P)$ represent the embeddings of the textual description and the protein sequence, respectively. Note that we follow the original implementation of the ProTrek Score  [^4] August 2024 version: https://www.biorxiv.org/content/10.1101/2024.08.01.606258v1.full.pdf as described in [8]. In contrast, the latest version of the paper defines the ProTrek Score with an additional division by the model temperature.
@@
EvoLlama $\text{EvoLlama Score}=\mathrm{sim}(\frac{1}{k}\sum_{i=1}^{k}\mathtt{Embed}(t),\frac{1}{k^{\prime}}\sum_{i=1}^{k^{\prime}}\mathtt{Embed}(t^{\prime}))$
@@
Keyword $\text{Keyword Recovery}=\frac{|K_{\text{pred}}\cap K_{\text{ref}}|}{|K_{\text{ref}}|}$
@@
Retrieval Accuracy [6] evaluates a model’s ability to correctly associate a textual function description $t$ with its corresponding protein sequence $P$. It is computed by comparing the similarity between the text description embedding and multiple protein embeddings, including one correct match and $T-1$ randomly retrieved negative samples. The metric is defined as the proportion of instances in which the correct protein sequence exhibits the highest similarity to the function description among all candidates. Note that the original implementation relies on ProteinCLAP [6] for both text and protein embeddings, whereas we adopt ProTrek [32] to achieve faster and more accurate representations, denoted as $\mathtt{Embed}$. The Retrieval Accuracy for a text-protein pair $(t,P_{1})$ is defined as:
Retrieval $\text{Retrieval Accuracy}@T=\mathbf{1}\left[\mathrm{sim}\left(\mathtt{Embed}(t),\mathtt{Embed}(P_{1})\right)=\max\limits_{j=1}^{T}\left\{\mathrm{sim}\left(\mathtt{Embed}(t),\mathtt{Embed}(P_{j})\right)\right\}\right]$

### v2

@@ 1 Introduction
Figure 1: (a) Visualization of proteins designed by our method, embedded using ESM C hayes2024simulating and projected with UMAP mcinnes2018umap . Random refers to proteins generated by randomly selecting amino acids according to their empirical distribution in SwissProt bairoch2000swiss . Random+ refers to proteins generated by selecting amino acids and incorporating fragments. Natural proteins randomly sampled from SwissProt (in gray) form a broad distribution, representing the diverse landscape of natural proteins. Proteins generated by Random sampling from all possible sequences (in red scatter points) cluster tightly at the periphery of the natural protein distribution, suggesting that random proteins are less diverse. Proteins generated by Random+ (in yellow scatter points) exhibit a much more diverse distribution than those generated by Random. (b) Performance on pLDDT ($\uparrow$). Our method improves pLDDT by 12% over Random+, exceeding the well-folded threshold. (c) Performance on PAE ($\downarrow$). Our method reduces PAE by 9% compared to Random+ and is the only model to surpass the well-folded threshold. Notably, it outperforms the state-of-the-art baseline model Pinal in both metrics.
@@ 2 Background
Proteins are macromolecules composed of linear chains of amino acids, with 20 standard amino acids serving as their fundamental building blocks. Continuous subsequences $S=\{s_{1},s_{2},\cdots,s_{m}\}$ within an amino acid sequence $P$, often referred to as fragments or segments, play crucial roles in determining the structure and function of proteins. In Figure 2, we utilize InterPro blum2025interpro , an integrated database and diagnostic tool, to identify these fragments within proteins. These fragments can be classified into 8 categories: Domain, Family, Homologous Superfamily, Repeat, Conserved Site, Active Site, Binding Site, and Post-Translational Modification (PTM). Details of fragment types are discussed in Appendix B. The fragments, along with their corresponding types and descriptions, are referred to as functional annotations, denoted by $\mathcal{A}_{\text{type}}(S)$ and $\mathcal{A}_{\text{desc}}(S)$.
@@ 3.2 Training
$\mathbf{W}_{\text{in}}$ $=\mathrm{Concat}(\mathbf{W}_{\text{tokens, in}},\mathbf{W}_{\text{fragments}})$ (1)
$\mathbf{H}_{\text{pre}}$ $=\mathrm{Concat}(\mathbf{H}_{t},\mathbf{W}_{\text{in}}[x_{<i}])$
@@
$\mathbf{W}_{\text{out}}$ $=\mathrm{Concat}(\mathbf{W}_{\text{tokens, out}},\mathbf{W}_{\text{fragments}}^{\mathrm{T}})$ (2)
$p(x_{i}=k|\mathbf{H}_{\text{pre}})$ $=\frac{\exp(\mathbf{H}_{\text{pre}}\mathbf{W}_{\text{out}}^{(k)})}{\sum\limits_{k^{\prime}\in V_{\text{tokens}}\cup S}\exp(\mathbf{H}_{\text{pre}}\mathbf{W}_{\text{out}}^{(k^{\prime})})}$
@@ 4.1 Experimental Setups
To comprehensively evaluate our approach, we consider the following metrics from four perspectives. Detailed definitions and implementations are provided in Appendix D. Sequence Plausibility. For rapid validation of designed protein sequences, we employ Perplexity (PPL) and Repetitiveness (Rep) welleck2019neural to evaluate sequence-level plausibility. Foldability. Following dai2024toward , we utilize ESMFold lin2022language to predict the 3D structures of designed proteins and compute the predicted Local Distance Difference Test (pLDDT) jumper2021highly as a per-residue measure of local confidence (0-100, higher is better). Additionally, we also compute the Predicted Aligned Error (PAE) jumper2021highly , which measures the confidence in the relative position of two residues within the predicted structure (lower is better). Consistent with previous researches jumper2021highly ; akdel2022structural ; varadi2022alphafold ; dai2024toward , we adopt pLDDT above 70 and PAE below 10 to define well-folded predicted structures. Language Alignment. To mitigate the high cost of wet-lab experiments, we employ oracle model-based metrics to evaluate the alignment between proteins and functions. Following dai2024toward , we use ProTrek su2024protrek , which evaluates alignment using cosine similarity, referred to as the ProTrek Score, within a joint embedding space. We also use EvoLlama liu2024evollama , fine-tuned on our downstream datasets, computing the EvoLlama Score as the cosine similarity between textual descriptions and predicted functions. Higher scores from both models indicate better alignment with the specified function. Additionally, we employ two retrieval-based metrics. The first is Keyword Recovery hayes2024simulating , which measures the recovery of function keywords in the designed proteins using InterProScan jones2014interproscan . And the second is Retrieval Accuracy liu2025text , which assesses how well the designed proteins aligns with their functional descriptions by evaluating whether their representations exhibit the highest similarity among all candidate pairs. Sequence Diversity. MMseqs2 steinegger2017mmseqs2 is used to compute sequence-level similarity between each pair of proteins within a batch, where all sequences are generated based on the same textual description. A higher score indicates a greater ability to design diverse protein sequences.
@@ 4.2 Designing Proteins from Function Keywords
Models
#Pairs
#Params
Sequence Plausibility Foldability Language Alignment (in %)
Sequence
Diversity
($\uparrow$)
PPL
($\downarrow$)
Rep
($\downarrow$)
pLDDT
($\uparrow$)
% > 70
($\uparrow$)
PAE
($\downarrow$)
% < 10
($\uparrow$)
ProTrek
Score ($\uparrow$)
Keyword
Recovery ($\uparrow$)
Retrieval
Accuracy ($\uparrow$)
Natural - 467.64 0.02 81.21 90.53 7.08 82.05 21.09 100.00 70.81 -
Random (U) - 2471.95 0.01 24.38 0.00 23.81 0.13 7.50 0.00 6.05 97.46
Random (E) - 3046.64 0.01 27.46 0.00 23.70 0.00 6.59 0.00 5.13 99.78
Random+ (E) - 966.24 0.01 62.38 32.65 17.23 9.28 3.29 0.00 5.79 98.97
ProteinDT 541K/729M 1405.70 0.11 38.70 0.20 26.25 0.00 3.89 0.05 7.43 99.72
ProteinDTFT 392K/729M 1860.43 0.04 38.66 1.04 23.90 0.42 6.28 1.08 16.57 99.32
Pinal 1.76B/2B 584.22 0.15 66.50 47.21 14.57 33.53 14.57 30.46 51.68 82.72
PAAG 130K/1.3B 2571.40 0.02 33.14 0.00 23.31 0.00 5.21 0.23 7.10 99.02
PAAGFT 392K/1.3B 2004.01 0.04 41.53 1.12 24.34 0.46 3.46 0.01 7.82 99.87
Chroma 45K/334M 1322.37 0.03 61.66 28.96 13.01 39.03 2.97 0.11 6.57 97.21
ESM3 539M/1.4B 279.78 0.33 59.79 31.49 17.40 21.37 3.76 5.49 11.97 96.77
ProDVa 392K/1.8B 656.04 0.01 75.88 77.00 6.39 83.88 14.43 30.34 44.77 98.58
@@ 4.3 Designing Proteins from Textual Descriptions
Models
#Pairs
#Params
Sequence Plausibility Foldability Language Alignment (in %)
Sequence
Diversity
($\uparrow$)
PPL
($\downarrow$)
Rep
($\downarrow$)
pLDDT
($\uparrow$)
% > 70
($\uparrow$)
PAE
($\downarrow$)
% < 10
($\uparrow$)
ProTrek
Score ($\uparrow$)
EvoLlama
Score ($\uparrow$)
Retrieval
Accuracy ($\uparrow$)
Natural - 318.15 0.02 80.64 81.27 9.20 65.73 27.00 60.33 84.85 -
Random (U) - 2484.03 0.01 22.96 0.16 24.85 0.56 1.03 36.23 6.89 97.01
Random (E) - 3136.88 0.01 25.77 0.20 24.71 0.60 1.04 34.11 6.78 99.56
Random+ (E) - 846.01 0.01 64.47 37.03 17.91 7.52 0.30 38.65 6.13 98.63
ProteinDT 541K/729M 1576.23 0.07 38.29 0.98 25.13 0.40 1.20 40.57 9.28 99.23
ProteinDTFT 712K/729M 1213.38 0.04 51.42 25.61 18.57 23.92 13.89 52.84 47.29 79.87
Pinal 1.76B/2B 308.97 0.13 75.25 68.97 10.96 58.44 17.50 53.42 57.95 82.96
PAAG 130K/1.3B 2782.70 0.02 28.39 0.07 25.38 0.10 1.29 34.39 7.06 99.15
PAAGFT 712K/1.3B 1332.35 0.04 50.37 23.86 19.96 21.99 10.04 49.69 33.66 86.09
Chroma 45K/334M 1370.21 0.03 59.18 20.17 15.03 28.62 2.10 40.10 7.33 96.13
ProDVa 712K/1.8B 415.63 0.02 76.86 76.35 8.66 68.06 17.40 51.10 59.07 83.29
@@ Results
(3) ProDVa demonstrates competitive sequence diversity compared to other baselines. We find that baselines that struggle to design well-folded or well-aligned proteins consistently exhibit higher sequence diversity, comparable to that of random methods. Furthermore, textual descriptions containing more constraints than simple keywords generally result in lower sequence diversity, suggesting that complex function descriptions may restrict the design space. By leveraging the top‑k sampling strategy during decoding, ProDVa achieves slightly better performance than Pinal on Mol‑Instructions and surpasses it by 15.86% on the CAMEO subset.
4.4 Unconditional Protein Generation
Table 3: Performance on the unconditional protein generation task (Best, Second Best). The fine-tuned models and ProDVa are initialized with the weights obtained in Section 4.3, without additional fine-tuning for this task.
Models
PPL
($\downarrow$)
Rep
($\downarrow$)
pLDDT
($\uparrow$)
% > 70
($\uparrow$)
PAE
($\downarrow$)
% < 10
($\uparrow$)
ProteinDTFT 593.06 17.92 47.79 0.02 26.56 0.00
PAAGFT 1327.98 3.55 50.32 23.83 19.95 22.24
Pinal 411.93 14.05 70.11 57.02 12.76 48.44
ProDVa 476.02 1.47 77.52 79.78 9.32 60.25
Although ProDVa is designed for conditional protein design tasks, we demonstrate that it can be adapted for unconditional protein generation. By fixing the input instruction to Design a novel protein sequence, the problem addressed by ProDVa effectively transforms into an unconditional protein generation task. To further ensure that the input acts as a no-op, we replace the retrieval method with the random selection of fragments from the training set described in Section 4.3. For a fair comparison, each model generates the same number of protein sequences as the size of the Mol-Instructions test set. The results are shown in Table 3, ProDVa outperforms all baseline models on the unconditional protein generation task. Compared to the state-of-the-art baseline, Pinal, ProDVa generates 22.76% and 11.81% more well-folded proteins in terms of pLDDT and PAE, respectively. Furthermore, compared to other fine-tuned models, ProDVa achieves substantially superior performance. These findings highlight the effectiveness of our approach in both conditional and unconditional protein generation settings.
@@ 6 Conclusion
Acknowledgement
The authors wish to thank all reviewers for their helpful comments and suggestions. We also thank Jie Wang for his insightful discussions. The corresponding authors are Yuanbin Wu, Tao Ji, Changzhi Sun and Man Lan. The computations in this research were performed using the CFFF platform of Fudan University.
@@ Appendix B Details of InterPro and Fragment Types
Table 4 presents the details of the 8 fragment types identified by InterPro.
Table 4: Descriptions of 8 fragment types identified by InterPro.
@@ C.1 Training Details
ProDVa is trained using the AdamW optimizer [53] with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a gradient clipping of 1.0. The maximum learning rate is set to $1\times 10^{-4}$, with a linear warmup over the first 5% of training steps. The learning rate remains constant after warmup and follows a "1-sqrt" decay schedule during the last 10% of training steps. The minibatch size is set to 4 to ensure memory efficiency, with an overall batch size of 64. During training, the weights of the Text Language Model are frozen. Training is conducted for 10K steps on the CAMEO subset and 20K steps on Mol-Instructions. The loss weights are set to $\alpha=\beta=0.2$.
@@ D.1 Sequence Plausibility Metrics
where $(x_{1},x_{2},\cdots,x_{n})$ is a sequence of tokens tokenized using the ProtGPT2 tokenizer. Additional PPL evaluations using alternative language models are provided in Appendix E.2.
@@ D.3 Language Alignment Metrics
The ProTrek Score is defined as the cosine similarity between a textual function description $t$ and a designed protein sequence $P$. Since ProDVa generates protein sequences directly, without an intermediate step for designing structure backbones, only the text and sequence encoders of ProTrek [32] are utilized. Specifically, we employ the pre-trained ProTrek 650M 333 https://huggingface.co/westlake-repl/ProTrek_650M_UniRef50 as the oracle model. The ProTrek Score is formally defined as:
$\text{ProTrek Score}=\mathrm{sim}(\tau_{t}(t),\tau_{p}(P))$
where $\mathrm{sim}(\cdot,\cdot)$ denotes the cosine similarity function, and $\tau_{t}(t)$ and $\tau_{p}(P)$ represent the embeddings of the textual description and the protein sequence, respectively. Note that we follow the original implementation of the ProTrek Score 444 August 2024 version: https://www.biorxiv.org/content/10.1101/2024.08.01.606258v1.full.pdf as described in [8]. In contrast, the latest version of the paper defines the ProTrek Score with an additional division by the model temperature.
@@
$\text{EvoLlama Score}=\mathrm{sim}(\frac{1}{k}\sum_{i=1}^{k}\mathtt{Embed}(t),\frac{1}{k^{\prime}}\sum_{i=1}^{k^{\prime}}\mathtt{Embed}(t^{\prime}))$
@@
$\text{Keyword Recovery}=\frac{|K_{\text{pred}}\cap K_{\text{ref}}|}{|K_{\text{ref}}|}$
@@
Retrieval Accuracy [6] evaluates a model’s ability to correctly associate a textual function description $t$ with its corresponding protein sequence $P$. It is computed by comparing the similarity between the text description embedding and multiple protein embeddings, including one correct match and T1T-1 randomly retrieved negative samples. The metric is defined as the proportion of instances in which the correct protein sequence exhibits the highest similarity to the function description among all candidates. Note that the original implementation relies on ProteinCLAP [6] for both text and protein embeddings, whereas we adopt ProTrek [32] to achieve faster and more accurate representations, denoted as $\mathtt{Embed}$. In our evaluation, $T$ is set to 20, representing the most challenging setting in [6]. The Retrieval Accuracy for a text-protein pair $(t,P_{1})$ is defined as:
$\text{Retrieval Accuracy}@T=\mathbf{1}\left[\mathrm{sim}\left(\mathtt{Embed}(t),\mathtt{Embed}(P_{1})\right)=\max\limits_{j=1}^{T}\left\{\mathrm{sim}\left(\mathtt{Embed}(t),\mathtt{Embed}(P_{j})\right)\right\}\right]$
@@
D.4 Sequence Diversity Metric
Sequence Diversity
Sequence Diversity measures a model’s ability to generate diverse protein sequences under identical conditions (i.e., given the same function as inputs). Given a textual description, the model is tasked with designing a batch of $N$ protein sequences, denoted as $\{P_{1},P_{2},\cdots,P_{N}\}$. Formally, Sequence Diversity is defined as:
$\text{Sequence Diversity}=\frac{2}{N(N-1)}\sum_{1\leq i\leq j\leq N}\mathrm{sim}(P_{i},P_{j})$
where MMseqs2 [35] is used to compute the sequence-level similarity, $\mathrm{sim}(\cdot,\cdot)$, between each pair of proteins $P_{i}$ and $P_{j}$.